OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 5 to 8. The perimeter of the larger rectangle is 72 inches. What is the perimeter of the smaller rectangle? A. 115.2 inches B. 45 inches C. 40 inches D. 16.4 inches

OpenStudy (anonymous):

It is either 'B' or 'C'

OpenStudy (swamijatin2):

solve this equation 5/8=72/x u will get the answer

OpenStudy (sepeario):

Very well said @swamijatin2

OpenStudy (swamijatin2):

thanx

OpenStudy (swamijatin2):

did u get the answer

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

115.2

OpenStudy (swamijatin2):

yaa right

OpenStudy (anonymous):

alright thank you

OpenStudy (swamijatin2):

your welcome

OpenStudy (anonymous):

it was 45

OpenStudy (anonymous):

so it was B

OpenStudy (swamijatin2):

how

OpenStudy (anonymous):

I think it was meant to be 5/8=x/72

OpenStudy (anonymous):

I solved it it was 5/8=x/72

OpenStudy (swamijatin2):

no it was 5/8=72/x x=72 x 8/5 x=115.2

OpenStudy (anonymous):

it is 5/8=x/72 360=8x 360/8=8x/8 45=x

OpenStudy (swamijatin2):

please wait

OpenStudy (swamijatin2):

you made me confused

OpenStudy (swamijatin2):

from my side i don't think it is 45